OpenStudy (anonymous):

find gcf of 18P^3 q^2,8pq^3,20pq^2

OpenStudy (anonymous):

Forgot this i wrote it wrong

OpenStudy (anonymous):

Oh nevermind

OpenStudy (anonymous):

I wrote it right

OpenStudy (anonymous):

first look at the numerical factors: 18=2*3*3 8=2*2*2 20=2*2*5 What set of factors do these three have in common? 2 Now do the same with the variable p. P*P*P p p So the greatest common factor is p. now for q. q*q q*q*q q*2 GCF is q*2. Now just put all the results together: 2*p*q*2 = 2pq^2